# Rigidity and Structural Asymmetry of Bounded Solutions

### 16 Mar 2026 v1

@@ Abstract
Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Volterra integral equation.
@@ 1 Main result
where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in L^{\infty}(\mathbb{C})$ where $L^{\infty}(\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$. As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
$\exists\rho\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<+\infty,$ (H)
The complex number $\rho$ is called the rotation number of $\eta$.
We are interested in the initial conditions of the bounded solutions of the previous differential equation. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
@@
Since $\eta\in L^{\infty}(\mathbb{C})$, the function
@@
For every $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
@@
For every $\eta\in L^{\infty}(\mathbb{C})$, let $\mu_{\eta}$ denote the function $\mu_{\eta}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
$\mu_{\eta}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
The function $\mu_{\eta}$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since $\eta\in L^{\infty}(\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
@@
Let $\eta\in L^{\infty}(\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
@@
where the function $\mu_{\eta}$ is defined in Notation 2. Furthermore, if $\eta$ satisfies the rotation number hypothesis (H), then
$\sup_{t\geq 1}\left|t\left(\psi_{\eta,w}\big(t\big)-\rho\frac{w-1}{w}t^{-i\Im(w)}\right)\right|<+\infty,$
where $\rho$ denotes the rotation number of $\eta$.
@@
By Hypothesis $\eta\in L^{\infty}(\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ almost everywhere, we obtain
@@
Now, prove the second item of the present Lemma. Suppose that the function $\eta$ satisfies the hypothesis (H). Then there exists $\tilde{c}>0$ such that
$\Big|\int_{1}^{t}\,\big(\eta(u)-\rho\big)du\Big|<\tilde{c},\quad\forall t\geq 1,$
where $\rho$ is the rotation number of $\eta$. Hence, the equation (4) can be written as
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w)t^{{\Re(w)}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,$
Using the integration by parts formula, we obtain
$\psi_{\eta,w}(t)=\rho\frac{w-1}{w}t^{-i\,\Im(w)}-(1-w^{2})t^{\Re(w)}\int_{t}^{+\infty}u^{-2-w}\int_{t}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{\eta,w}(t)-\rho\frac{w-1}{w}t^{-i\,\Im(w)}\right|<2\tilde{c}\frac{|1-w^{2}|}{1+\Re(w)}t^{-1}.$
@@
Let $\eta\in L^{\infty}(\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the function $\mu_{\eta}$ satisfies:
$\forall\beta\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,\beta)\neq 0.$
@@ 2 Main proposition
Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
@@
For every $s:=\sigma+i\tau\in I$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following integral equation:
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
@@
3 Proof of the Theorem 4
Lemma 7.
For every $s\in I$ with $\Re(s)=\sigma$, the function $\delta_{\eta,s}$, defined in Notation (5), satisfies the following Volterra integral equation:
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
Proof.
By the Proposition 6 every $s\in I$, we have
$+\int_{1}^{t}u^{-\frac{3}{2}-i\,\tau}\,\eta(u)\ du.$
By definition of the set II in Notation 5, we have in particular,
$t^{-\frac{1}{2}}\delta_{\eta,1+i\Im(s)}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du$
@@ 3 Proof of the Theorem 4
$\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\Im(s)}(v)\,dv\,du.$
@@
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\Im(s)}(t)\Big)=\sigma(\sigma-1)t^{-2}\delta_{\eta,s}(t),$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\Im(s)}(1)$ and $\delta_{\eta,s}^{\prime}(1)=\delta_{\eta,1+i\Im(s)}^{\prime}(1)$. Solving this Cauchy problem gives the explicit representation
$\delta_{\eta,s}(t)=\delta_{\eta,1+i\Im(s)}(t)-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
@@
Let $s\in B$ and $\sigma=\Re(s)$. By the Lemma 7, we have
$\delta_{\eta,s}(t)$ $=\delta_{\eta,1+i\Im(s)}(t)$ (8)
$-\frac{\sigma(1-\sigma)}{2\sigma-1}\int_{1}^{t}\left[\left(\frac{t}{v}\right)^{\sigma}-\left(\frac{t}{v}\right)^{1-\sigma}\right]v^{-1}\delta_{\eta,1+i\Im(s)}(v)\,dv.$
By equation (3), we have
$\delta_{\eta,1+i,\Im(s)}(t)$ $=t\Big[1-i\ \Im(s)\int_{1}^{t}u^{-2-i\Im(s)}\eta(u)\,du\Big]$
$-1-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\eta(u)\,du,\quad\forall t\geq 1,$
which can be written as
$\delta_{\eta,1+i\Im(s)}(t)$ $=-t\mu\big(1+i\Im(s)\big)+i\ \Im(s)t\int_{t}^{+\infty}u^{-2-i\Im(s)}\eta(u)\,du$
$-1-\rho\frac{1-i\ \Im(s)}{i\Im(s)}\big(1-t^{-i\Im(s)}\big)$
$-(1-i\ \Im(s))\int_{1}^{t}u^{-1-i\Im(s)}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Applying the integration by parts formula to the last integral of the previous equation and using the hypothesis (H), there exists $M_{\Im(s)}>0$ such that
$\Big|\delta_{\eta,1+i\Im(s)}(t)+t\mu\big(1+i\Im(s)\big)\Big|<M_{\Im(s)},\quad\forall t>1.$ (9)
By hypothesis, $\mu_{\eta}(1+i\beta)\neq 0$. Assume $\Im(\mu_{\eta}(1+i\beta))\neq 0$ (the real case is analogous). Equation (9) implies that for sufficiently large $t$, $\Im(\delta_{\eta,1+i\Im(s)}(t))$ maintains a constant sign and grows linearly. It follows from (8) that:
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}\to 0^{+}\text{ as }t\to+\infty.$ (10)
In fact
$\frac{\Im\big(\delta_{\eta,s}(t)\big)}{\Im\big(\delta_{\eta,1+i\Im(s)}(t)\big)}=1-\frac{\sigma(1-\sigma)}{2\sigma-1}I(t),$

### 15 Apr 2026 v3

@@ Abstract
Keywords: Non-homogeneous complex linear differential equation, bounded solutions, Euler differential equation, Volterra integral equation.
@@ 1 Introduction
where $w\in\mathbb{C}$ such that $\Im(w)\neq 0$ is the parameter and the non-homogeneous term $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ where $\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ is the set of bounded and locally integrable functions from $1,+\infty)$ to $\mathbb{C}$. We are interested in the initial conditions of the bounded solutions of the previous differential equation. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
@@
Since $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, the function
@@
For every $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}$, we denote by $\psi_{\eta,w}(\cdot)$ the unique continuous solution of the differential equation (2) given by equation (3).
@@
For every $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, let $\mu_{f}$ denote the function $\mu_{f}:\mathbb{C}_{+}\to\mathbb{C}$, defined as
$\mu_{f}(w)=-1-(1-w)\int_{1}^{+\infty}u^{-1-w}f(u)\,du,\quad\forall w\in\mathbb{C}_{+}.$
The function $\mu_{f}$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since $f\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
@@
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ and $w\in\mathbb{C}_{+}$. Consider the continuous solution $\psi_{\eta,w}(\cdot)$ of the differential equation (2), as introduced in Notation 1. Then,
@@
where the function $\mu_{\eta}$ is defined in Notation 2.
@@
By Hypothesis $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ then there exists a constant $c>0$ such that $|\eta(t)|<c$ for all $t\geq 1$, we obtain
@@
2 Main result
As formulated in the Dynamical Conjecture in [1], we assume that the function $\eta$ satisfies the following Rotation number hypothesis:
$\exists\rho_{\eta}\in\mathbb{C}^{*}:\quad\sup_{t\geq 1}\Big|\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du\Big|<+\infty,$ (H)
The complex number $\rho_{\eta}$ is called the rotation number of $\eta$. Notice that if $\rho_{\eta}$ exists, then it is unique; this can be proved by contradiction. Denote by $\mathcal{L}[\eta]$ the following transformation
$\mathcal{L}[\eta](t):=\int_{1}^{t}\,\big(\eta(u)-\rho_{\eta}\big)du,\quad\forall t\geq 1.$
The Rotation number hypothesis implies that $\mathcal{L}[\eta]\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. Hence, the function $\mu_{\mathcal{L}[\eta]}$ is defined on $\mathbb{C}_{+}$ as given in the Notation 2.
@@ 2 Main result
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). Suppose that the functions $\mu_{\eta}$ and $\mathcal{L}[\eta]$ satisfies:
$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\mathcal{L}[\eta]}(1+i\,x)\neq\frac{\rho_{\eta}-ix}{1+ix}-\frac{1}{|1+i\,x|^{2}}.$
@@
Remark 5.
In the proof of the Lemma 9, we note that
$\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big]=\lim_{r\to_{>}0}\mu_{\eta}(r+i\,x),$
which is the analytic extension of the function $\mu_{\eta}$ to the line $\Re(z)=0$. Consequently, the condition of the previous theorem is equivalent to
$\forall x\in\mathbb{R}^{*}:\quad\mu_{\eta}(1+i\,x)\neq 0\quad\text{and}\quad\mu_{\eta}(i\,x)\neq 0.$
@@ 3 Main proposition
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. Consider the continuous solutions of the differential equation (2) as given in Notation 1. For every $w\in I$, we denote by $\delta_{w}$ the function
@@
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Volterra integral equation:
$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
Lemma 8.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$. For every $s:=\sigma+i\tau\in I$, the functions $\delta_{\eta,s}$ and $\delta_{\eta,1+i\tau}$, defined in Notation (6), satisfies the following Euler differential equation:
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),\quad\forall t\geq 1,$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$.
@@
By definition of the set II in Notation 6, we have in particular,
$t^{-\frac{1}{2}}\delta_{\eta,1+i\tau}(t)=$ $\frac{1}{4}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du$
@@
$\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,s}(v)\,dv\,du$
$-\frac{1}{4}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{\eta,1+i\tau}(v)\,dv\,du.$
@@
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
with initial conditions $\delta_{\eta,s}(1)=\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$. ∎
Proof of the Proposition 7.
By the Lemma 8 every $s:=\sigma+i\,\tau\in I$, we have
$\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-2}\delta_{\eta,s}(t),$
which can be written as
$t^{1-\sigma}\frac{d^{2}}{dt^{2}}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\Big)=-\sigma(1-\sigma)t^{-1-\sigma}\delta_{\eta,s}(t),$
Use the fact that $\delta_{\eta,s}(1)-\delta_{\eta,1+i\tau}(1)=0$ and $\dot{\delta}_{\eta,s}(1)-\dot{\delta}_{\eta,1+i\tau}(1)=0$ and use the integration by part formula,
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)\int_{1}^{t}\,u^{-\sigma}\frac{d}{du}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
$-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\big(\delta_{\eta,s}(u)-\delta_{\eta,1+i\tau}(u)\big)\,du$
$=-\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,s}(u)\,du,$
We obtained
$t^{1-\sigma}\frac{d}{dt}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$ $-(1-\sigma)t^{-\sigma}\big(\delta_{\eta,s}(t)-\delta_{\eta,1+i\tau}(t)\big)$
$+\sigma(1-\sigma)\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
$=0,\quad\forall t\geq 1.$
That can be written as
$\frac{d}{dt}\Big(t^{-1+\sigma}\big(\delta_{\eta,1+i\tau}(t)-\delta_{\eta,s}(t)\big)\Big)$ $=\sigma(1-\sigma)t^{-2(1-\sigma)}\int_{1}^{t}\,u^{-1-\sigma}\delta_{\eta,1+i\tau}(u)du$
$+\frac{d}{dt}\big(t^{-1+\sigma}\delta_{\eta,s}(t)\big),\quad\forall t\geq 1,$
Integrate
$\delta_{\eta,1+i\tau}(t)$ $=\sigma(1-\sigma)t^{1-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\delta_{\eta,1+i\tau}(v)\,dv\ du$
$+\delta_{\eta,s}(t),\quad\forall t\geq 1,$
4 Proof of the Theorem 4
Lemma 9.
Let $\eta\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$ satisfy the Rotation Number Hypothesis (H). For every $x\in\mathbb{R}^{*}$, the function $\delta_{\eta,1+ix}$, defined in Notation (6), satisfies the following equation
$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
where
$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
Proof.
Let $x\in\mathbb{R}^{*}$. By the Notation 6 and equation (3), we have
$\delta_{\eta,1+i,x}(t)$ $=t\Big[1-i\,x\int_{1}^{t}u^{-2-ix}\eta(u)\,du\Big]$
$-1-(1-i\,x)\int_{1}^{t}u^{-1-ix}\eta(u)\,du,\quad\forall t\geq 1,$
which can be written as
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)+i\,xt\int_{t}^{+\infty}u^{-2-ix}\eta(u)\,du$
$-1-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)$
$-(1-i\,x)\int_{1}^{t}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Use the Rotation number hypothesis (H) and use the integration by formula
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
$-1-|1+i\,x|^{2}\int_{1}^{+\infty}u^{-2-ix}\int_{1}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du$
$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
Equivalent to,
$\delta_{\eta,1+ix}(t)$ $=-t\mu_{\eta}\big(1+ix\big)-\rho\frac{1-i\,x}{ix}\big(1-t^{-ix}\big)+\rho\frac{i\,x}{1+ix}t^{-ix}$
$-1-|1+i\,x|^{2}\frac{1}{ix}\big(1+\mu_{\mathcal{L}[\eta]}(1+i\,x)\big)$
$+i\,xt\int_{t}^{+\infty}u^{-2-ix}\big(\eta(u)-\rho\big)\,du$
$+(1-i\,x)\int_{t}^{+\infty}u^{-1-ix}\big(\eta(u)-\rho\big)\,du,\quad\forall t\geq 1.$
or,
$\delta_{\eta,1+ix}(t)$ $=-t\alpha_{x}+\beta_{x}$
$-\rho\frac{1}{ix(1+ix)}t^{-ix}$
$+i\,x(2+ix)t\int_{t}^{+\infty}u^{-3-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,\,du$
$+|1+i\,x|^{2}\int_{t}^{+\infty}u^{-2-ix}\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\,du,\quad\forall t\geq 1.$
where
$\alpha_{x}:=\mu_{\eta}\big(1+ix\big)\ \text{and}\ \beta_{x}:=\frac{|1+i\,x|^{2}}{ix}\Big[\frac{\rho_{\eta}-ix}{1+ix}-\mu_{\mathcal{L}[\eta]}(1+i\,x)-\frac{1}{|1+i\,x|^{2}}\Big].$
Using the hypothesis (H), we have
$\sup_{u\geq t\geq 1}\Big|\int_{t}^{u}\,\big(\eta(v)-\rho\big)\,dv\Big|<+\infty,$
Then
$\sup_{t\geq 1}\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+ix}(u)\,du+\alpha_{x}t-\beta_{x}\ln(t)\Big|<+\infty,$
@@ 4 Proof of the Theorem 4
Let $s:=\sigma+i\,\tau\in B$. By the Lemma 9, there exists $M_{\tau}>0$ such that
$\Big|\int_{1}^{t}\,u^{-1}\delta_{\eta,1+i\tau}(u)\,du+\alpha_{\tau}t-\beta_{\tau}\ln(t)\Big|<M_{\tau},\quad\forall t\geq 1,$ (8)
where by the hypothesis of the present Theorem we have $\alpha_{\tau}\neq 0$ and $\beta_{\tau}\neq 0$. Then there exists $\omega_{\tau}\in\mathbb{C}^{*}$ such that
$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0\quad\text{or}\quad\Im\big(\omega_{\tau}\alpha_{\tau})\Im\big(\omega_{\tau}\beta_{\tau})>0.$
Suppose that
$\Re\big(\omega_{\tau}\alpha_{\tau})\Re\big(\omega_{\tau}\beta_{\tau})>0.$ (9)
The proof is analogous for the imaginary part. Fix $\omega_{\tau}$ satisfying the previous equation and define
$\tilde{\delta}_{1+i\tau}(t):=\Re\big(\omega_{\tau}\phi_{\eta,\tau}(t)\big)\quad\text{and}\quad\tilde{\delta}(t):=\Re\big(\omega_{\tau}\delta_{\eta,s}(t)\big).$ (10)
Equation (8) implies
$\Big|\int_{1}^{t}\,u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du+\Re\big(\omega_{\tau}\alpha_{\tau})t-\Re\big(\omega_{\tau}\beta_{\tau})\,\ln(t)\Big|<M_{\tau}\,|\omega_{\tau}|,\quad\forall t\geq 1,$ (11)
For every $c>1$, denote
$\mathcal{A}_{c}(t):=\frac{1}{2}+\frac{1}{\pi}\arctan(t-c);\quad\forall t\geq 1.$
For every $c>1$ and $g\in\mathcal{B}_{\mathrm{loc}}([1,\infty),\mathbb{C})$, consider the following transformation
$\Phi_{c}[g](t):=c\,t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}g(u)\,du,\quad\forall t\geq 1.$
Since $\lim_{t\to+\infty}\mathcal{A}_{c}(t)=1$ for every $c>1$, by equation (11), we have
$\forall c>1:\quad\lim_{t\to+\infty}\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\big|\Re\big(\omega_{\tau}\alpha_{\tau})\big|>0.$ (12)
Using equation (9), there exist a sufficiently large $c_{s}>1$ and $t_{c_{s}}>1$ such that for all $c>c_{s}$ and $t>t_{c_{s}}$ we have
$\Big|t^{-1}\int_{1}^{t}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|=\max_{z\in[1,t]}\Big|z^{-1}\int_{1}^{z}\mathcal{A}_{c}(u)u^{-1}\tilde{\delta}_{1+i\tau}(u)\,du\Big|,$ (13)
Implies
$\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|=\max_{z\in[1,t]}\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\Big|,\quad\forall c>c_{s},\quad\forall t>t_{c_{s}}.$ (14)
The purpose of introducing $\mathcal{A}_{c}$ is to obtain the previous equality. It ensures that the peaks of the function $t\mapsto\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|$ are bounded above by the limiting value which is equal to $c|\Re\big(\omega_{\tau}\alpha_{\tau})|$.
By the Proposition 7, for all $t\geq 1$ we have
$\frac{t^{-1}}{\sigma(1-\sigma)}\Big(\tilde{\delta}_{1+i\tau}(t)$ $-\tilde{\delta}_{s}(t)\Big)$
$=t^{-\sigma}\int_{1}^{t}\,u^{-2(1-\sigma)}\int_{1}^{u}\,v^{-1-\sigma}\tilde{\delta}_{1+i\tau}(v)\,dv\ du,$
Applying the transformation $\Phi_{c}$,
$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
$=c\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}\,\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\mathcal{A}_{c}(z)z^{-1}\tilde{\delta}_{1+i\tau}(z)\,dz\,dv\ du,$
Use the integration by part formula,
$\frac{t}{\sigma(1-\sigma)}\Big(\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)-\Phi_{c}[\tilde{\delta}_{s}](t)\Big)$
$=\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,\frac{v^{-1+\sigma}}{\mathcal{A}_{c}(v)}\Phi_{c}[\tilde{\delta}_{1+i\tau}](v)\,dv\,du$
$-\int_{1}^{t}\mathcal{A}_{c}(u)u^{-\sigma}\int_{1}^{u}\,v^{-2(1-\sigma)}\int_{1}^{v}z\Big(\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}\Big)\,\Phi_{c}[\tilde{\delta}_{1+i\tau}](z)\,dz\,dv\,du,$
We have
$\frac{4}{\pi}\frac{d}{dz}\frac{z^{-\sigma}}{\mathcal{A}_{c}(z)}=-\frac{\sigma\,z^{-\sigma-1}\left(\frac{\pi}{2}+\arctan(z-c)\right)+\frac{z^{-\sigma}}{1+(z-c)^{2}}}{\left(\frac{\pi}{2}+\arctan(z-c)\right)^{2}}<0,\quad\forall z\geq 1.$
By equation (14), for all $c>c_{s}$ and for all $t\geq t_{c_{s}}$ , we obtain
$t\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|<t\Big|\Phi_{c}[\tilde{\delta}_{s}](t)\Big|+\sigma(1-\sigma)\Big|\Phi_{c}[\tilde{\delta}_{1+i\tau}](t)\Big|K(\sigma,c,t),$ (15)